OpenStudy (anonymous):

The sine of 12 degrees is equal to 0.208, rounded to the nearest thousandth. What is the measure of an angle whose cosine is also equal to 0.208?

geerky42 (geerky42):

Hint: \[\large \cos\left(90^o-x\right) = \sin(x)\]

OpenStudy (anonymous):

is it 78 ?

geerky42 (geerky42):

yep